During a lab experiment, 3.82 g magnesium nitride and 7.73 g water are placed in a reaction flask.

At the end of the experiment, the amount of magnesium oxide produced is 3.60 g.

What is the percent yield of the reaction?
1
Expert's answer
2020-06-11T10:28:11-0400

Mg3N2+6H2O= 3Mg(OH)2+2NH3

3.6/40 = 0.09.

3.72 / 100 = 0.0372.

100 x 0.09 /0.0372 x 3= 80,6451613%.